# HunyuanImage 3.0 Technical Report

### v1

@@ 2.3 Reasoning Dataset Construction
Text to Text. To enhance the mdoel’s instruction following and logical-reasoning ability, we curated a diverse textual corpus of real-world image generation prompts. This corpus spans photorealistic rendering, artistic and stylistic renderings, UI and poster design tasks, knowledge-driven queries, and scientific or technical visualizations. By covering a broad spectrum of user intents, domains, and complexity levels, the model trained with T2T data can parse nuanced requirements, resolve ambiguities, and produce coherent, stepwise textual reasoning that faithfully maps instructions to precise image captions.
@@ 4.2 Post-training
The post-training optimization of our model is a multi-stage process designed to systematically refine its generative capabilities. We first conduct SFT on a meticulously curated dataset of human-annotated examples. Following this, DPO [39] is implemented to effectively address and reduce physical distortions. We then utilize MixGRPO [40] to enhance the critical aspects of text-image alignment, realism, and aesthetic appeal. The final refinement is achieved through the application of SRPO [41] and a novel, in-house Reward Distribution Alignment (ReDA) method, which together are instrumental in further elevating the realism and clarity of the generated images.
@@
For the SFT stage, we collect carefully curated high quality images encompassing diverse categories, including landscapes, portraits, animals, OCR, and others. We implement a multi-stage training strategy where subsequent stages incorporate progressively higher-quality training samples.
@@
We utilize DPO to address common issues of structural deficiency in image generation. The training data is prepared by first generating a corpus of images from the SFT model. These images are subsequently annotated to create a paired dataset of high-quality and low-quality samples. This dataset serves as a preference signal, which is applied to effectively suppress distortions and improve visual appeal.
@@
MixGRPO is an efficient online reinforcement learning framework that extends GRPO to flow-based models through a hybrid ODE–SDE sampling strategy. We apply MixGRPO with both open-source and proprietary reward models to optimize aesthetics (style, composition, lighting), mitigate distortions, and reduce artifacts. In addition, we refine advantage estimation to accelerate convergence, and demonstrate the scalability of MixGRPO to large-scale training regimes, achieving stronger alignment with human preferences.
@@
We’ve developed a novel reward distritution alignment algorithm, ReDA, for the post-training phase of our generative model. ReDA effectively improves visual quality by minimizing the divergence between the model’s generated outputs and a high-reward distribution, which is defined by a diverse set of high-quality images from various genres.
@@ 7 Project Contributors
  • Project Sponsors: Jie Jiang, Linus, Peng Chen, Yuhong Liu
@@
    • Algorithm & Pretraining: Miles Yang, Fanbin Lu, Jian-Wei Zhang, Shi-Xue Zhang, Zijian Zhang
    • Post Training: Lucas Wang, Chunyu Wang, Hangting Chen, Hao Wen, Junzhe Li, Lucaz Liu, Tao Zhang, Xiangwei Shen, Yingfang Zhang, Ying Dong, Yixuan Shi, Yutao Cui, Zheng Yuan, Zhengkai Jiang, Zhimin Li
  • Contributors (Listed alphabetically): Bing Wu, Chao Zhang, Donghao Li, Fang Yang, Jiale Tao, Jianchen Zhu, Jiaxin Lin, Jingmiao Yu, Kai Wang, Kipper Gong, Lei Wang, Linqing Wang, Pengfei Wan, Penghao Zhao, Qi Tian, Qinglin Lu, Qixun Wang, Senhao Xie, Shu Liu, Siyu Cao, Tiankai Hang, Tianpeng Gu, Weigang Zhang, Weijie Kong, Weiyan Wang, Xiaofeng Yang, Xiusen Gu, Xuan Yang, Yangyu Tao, Yanxin Long, Yepeng Zhang, Yiji Cheng, Yu Liu, Yuanbo Peng, Yue Wu, Yuyang Peng, Zhantao Yang, Zhenxi Li, Zhiyuan Zhao, Zihao Zhang

### v2

@@ 2.3 Reasoning Dataset Construction
Text to Text. To enhance the model’s instruction following and logical-reasoning ability, we curated a diverse textual corpus of real-world image generation prompts. This corpus spans photorealistic rendering, artistic and stylistic renderings, UI and poster design tasks, knowledge-driven queries, and scientific or technical visualizations. By covering a broad spectrum of user intents, domains, and complexity levels, the model trained with T2T data can parse nuanced requirements, resolve ambiguities, and produce coherent, stepwise textual reasoning that faithfully maps instructions to precise image captions.
@@
Text&Image to Text&Image. To enhance the model’s capability for complex image editing, we constructed the TI2TI corpus, a diverse dataset of editing pairs. The corpus consists of source images, complex editing instructions, and the corresponding ground-truth edited images. Crucially, for each instruction, we annotate a detailed editing trace. This trace deconstructs the user’s complex instruction into a sequence of atomic operations. This collection of source images, instructions, editing traces, and final images is designed to train the model to perform step-by-step reasoning for image editing, thereby significantly improving its ability to follow complex and compositional user instructions.
@@ 4.2 Post-training
The post-training optimization of our model is a multi-stage process designed to systematically refine its generative capabilities. We first conduct SFT on a meticulously curated dataset of human-annotated examples. Following this, DPO [39] is implemented to effectively reduce visual artifacts and physical distortions. We then utilize MixGRPO [40] to enhance the critical aspects of text-image alignment, realism, and aesthetic appeal. The final refinement is achieved through the application of SRPO [41] and a novel, in-house Reward Distribution Alignment (ReDA) method, which together are instrumental in further elevating the realism and clarity of the generated images.
@@
For the SFT stage, we curate a high-quality text-to-image dataset spanning diverse domains, including landscapes, portraits, and OCR. This is further expanded into a comprehensive editing dataset by systematically pairing image types with specific editing operations, governed by stringent filters to ensure spatial consistency and identity preservation. To enhance the instruction-following capabilities of the model, we integrate sophisticated reasoning data. Finally, a multi-stage training strategy is employed, where subsequent phases progressively incorporate higher-fidelity samples to refine performance.
@@
We employ DPO to mitigate prevalent issues in text-to-image and editing tasks, such as structural distortion, sampling instability, and perceptible synthetic artifacts. Our DPO is structured into two distinct stages. In the first stage, we utilize an extensive corpus of paired data to bolster sampling stability and ensure consistency in editing. The second stage focuses on refining visual realism by leveraging a curated, high-quality subset of samples. This hierarchical training strategy effectively suppresses structural defects while significantly reducing visual artifacts and enhancing the overall aesthetic fidelity of the generated images.
@@
MixGRPO is an efficient online reinforcement learning framework that extends GRPO to flow-based models through a hybrid ODE–SDE sampling strategy. We apply MixGRPO with proprietary reward models to optimize aesthetics (style, composition, lighting), mitigate distortions, and reduce artifacts. In addition, we refine advantage estimation to accelerate convergence, and demonstrate the scalability of MixGRPO to large-scale training regimes, achieving stronger alignment with human preferences. To address a variety of image-to-image tasks, including image editing and identity (ID) preservation, we employ a balanced multi-task joint training strategy. This approach involves iteratively training multiple reward models, each targeting a different quality dimension. For the face ID preservation task, we specifically investigated and tailored a suitable reward model. Furthermore, we curated a challenging and diverse dataset of image-to-image pairs to guide model convergence more efficiently. Through these comprehensive explorations, we achieved significant improvements in the model’s ability to preserve non-edited regions, maintain facial identity, and enhance overall image quality.
@@
Reward Distribution Alignment optimizes the post-training phase by minimizing the divergence from a high-reward prior. We employ task-specific projectors to map generations into a compressed space, allowing for targeted optimization of metrics like identity consistency and realism. When reference images are provided, we introduce a transition-based objective: instead of comparing static features, we calculate the vector difference (transition) between the reference and the generated images. This formulation significantly boosts training efficiency. Finally, the algorithm is designed to be data-centric, effectively leveraging improvements in dataset scale and quality.
4.3 Distillation
To accelerate generation speed, we propose a novel distillation framework that scales MeanFlow [42] for distilling the 80B-parameter unified multimodal model. Within this framework, we effectively mitigate the training instability issue, and further incorporate trajectory distribution alignment into the MeanFlow objective to enhance the few-step generation performance. Ultimately, we reduce the Number of Function Evaluations (NFE) to 4–8 while preserving competitive model performance.
@@ 7 Project Contributors
  • Project Sponsors: Jie Jiang, Liefeng Bo, Peng Chen, Yuhong Liu
@@
    • Algorithm & Pretraining: Miles Yang, Fanbin Lu, Jian-Wei Zhang, Qixun Wang, Shi-Xue Zhang, Yiji Cheng, Zijian Zhang
    • Post Training: Lucas Wang, Chunyu Wang, Hangting Chen, Hao Wen, Junzhe Li, Lucaz Liu, Ringo Ling, Tao Zhang, Xiangwei Shen, Xindi Yang, Yingfang Zhang, Ying Dong, Yixuan Shi, Yutao Cui, Zheng Yuan, Zhengkai Jiang, Zheyuan Liu, Zhimin Li
  • Contributors (Listed alphabetically): Bing Wu, Chao Zhang, Donghao Li, Fang Yang, Jiale Tao, Jianchen Zhu, Jiaxin Lin, Jingmiao Yu, Kai Wang, Kipper Gong, Lei Wang, Linqing Wang, Pengfei Wan, Penghao Zhao, Qi Tian, Qinglin Lu, Senhao Xie, Shu Liu, Shuang Chen, Siyu Cao, Tiankai Hang, Tianpeng Gu, Weigang Zhang, Weijie Kong, Weiyan Wang, Xiaofeng Yang, Xiusen Gu, Xuan Yang, Yangyu Tao, Yanxin Long, Yepeng Zhang, Yu Liu, Yuanbo Peng, Yue Wu, Yuyang Peng, Zhantao Yang, Zhenxi Li, Zhiyuan Zhao, Zihao Zhang